OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8

OpenStudy (anonymous):

do u know how to do it

OpenStudy (jdoe0001):

can you solve the 1st equation for "y"?

OpenStudy (anonymous):

yeeah i guess

OpenStudy (jdoe0001):

so, what do you end up?

OpenStudy (anonymous):

y=-4x+11

OpenStudy (jdoe0001):

or y = 11-4x... right

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

\(\bf 4x + y = 11\\ x + 2y = 8\\ --------------\\ y = 11-4x\\ \textit{now we } \color{blue}{\textit{substitute }} \textit{"y" in the 2nd equation}\\ \textit{since y = 11-4x, then we could say that}\\ x + 2y = 8 \implies x + 2(11-4x) = 8\) if you were to solve THAT for "x", what would that give you?

OpenStudy (anonymous):

-7x+22=8

OpenStudy (anonymous):

-7x=-12

OpenStudy (jdoe0001):

8 -22 = -12?

OpenStudy (anonymous):

-14

OpenStudy (jdoe0001):

-7x = -14 => x = 2 now that we know that "x = 2" we can plug it on either equation to find out what "y" is, by solving for "y"

OpenStudy (jdoe0001):

\(\bf 4x + y = 11\\ x + 2y = 8\\ --------------\\ 4(2) + y = 11\)

OpenStudy (jdoe0001):

8 + y = 11 => y = 3

OpenStudy (jdoe0001):

let's see if that's the case, x = 2, y = 3 \(\bf 4x + y = 11\\ x + 2y = 8\\ --------------\\ 4(2) + 3 = 11\\ 2 + 2(3) = 8\)